OpenStudy (anonymous):

Naomi has a neon tetra fish and a guinea pig the fish is 25 meters long the Guinea pig is 25 centimeters long what is the difference between the length of the fish and the length of a guinea pig

OpenStudy (danjs):

a 25 meter long tetra

OpenStudy (danjs):

1 cm = 10^(-2) m large differnce 25cm = 25*10^(-2) m or 0.25m

OpenStudy (danjs):

25 divided by 0.25 is 100 times larger

OpenStudy (mathmale):

Wow...a tetra fish about 76 feet long? I'm getting out of the water. Seriously, what is 25 meters less 25 centimeters, if 100 cm = 1 m?

OpenStudy (vickyisthesmartone123):

100*25=2500

OpenStudy (anonymous):

whats the answer